OpenStudy (anonymous):

integral of dx/2x^2 +4x +3

ganeshie8 (ganeshie8):

complete the square in the denominator and try a trig substitution

OpenStudy (anonymous):

well... could it be like this... \[\int\limits_{}^{}\frac{ dx }{ 2x ^{2}+4x+3 }= \frac{ 1 }{ 2 }\int\limits_{}^{}\frac{ dx }{ (x+1)^{2}+1}\]

ganeshie8 (ganeshie8):

Excellent ! looks like u substitution will do : u = x+1

OpenStudy (thomas5267):

I am often amazed that how dumb I am. I was going to do a trig substitution lol.

OpenStudy (anonymous):

so then,\[(\tan ^{-1} x+1) +c\]

ganeshie8 (ganeshie8):

haha i was about to do the same lol

ganeshie8 (ganeshie8):

looks good^^ don't forget the 1/2 which you have factored out earlier

OpenStudy (anonymous):

ah...haha i missed it